# Quantum Dynamic Programming

### v1

@@ IV.2.1 Distance bounds after a DME implementation
$\frac{1}{2}\lVert\rho^{\prime}_{n}-\tilde{\rho}_{n}\rVert_{1}\geq\frac{1}{2}\left\lVert(x^{\prime}_{n}-x_{n})\psi_{n}+(x_{n}-x^{\prime}_{n})\frac{\mathmybb{1}_{\rm rel}}{2}\right\rVert_{1}=\frac{\lvert x_{n}-x^{\prime}_{n}\rvert}{2}.$ (109)
@@
$\frac{1}{2}\lVert\rho^{\prime}_{n}-\tilde{\rho}_{n}\rVert_{1}\leq\frac{1}{2}\left\lVert\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})-\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})\right\rVert_{\rm 1}\leq\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}\right\rVert_{\rm Tr}<\frac{7L^{2}}{M}\ .$ (110)
Therefore,
$0<x^{\prime}_{n}<x_{n-1}+\frac{7L^{2}}{M}<\frac{L\pi+2}{4L\pi}\epsilon\ ,$ (111)
since we set $\frac{7L^{2}}{M}<\frac{L\pi-2}{4L\pi}\epsilon$ in Eq. (108).
In the next subsection, we describe a subroutine to reduce the mixedness factor $x^{\prime}_{n}$ into $x_{n}\leq\frac{\epsilon}{L\pi}$ so that we can continue the DME implementation without an accumulating error. However, if we eventually change the factor $x^{\prime}_{n}$, it is more important to get a bound for $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}$ corresponding to the associated pure state, rather than $\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}$ as in Lemma 14.
@@
The resulting state $\rho^{\prime}_{n}=(1-x_{n}^{\prime})\psi_{n}+x^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}$ of Lemma 14 has the associated pure state, such that
$\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}$ $\leq\delta_{n}\ ,$ (112)
where $\delta_{n}$ is defined in Eq. (91).
@@
The distance in Eq. (112) follows the triangle inequality $\lVert\tau-\psi_{n}\rVert_{1}\leq\lVert\tau-\rho^{\prime}_{n}\rVert_{1}+\lVert\rho^{\prime}_{n}-\psi_{n}\rVert_{1}$. The first term in the RHS is already bounded by $\delta_{n}-\frac{\epsilon}{4}$ in Lemma 14. The second term $\frac{1}{2}\lVert\rho^{\prime}_{n}-\psi_{n}\rVert_{1}=\frac{x^{\prime}_{n}}{2}$ by definition. We then use Eq. (111) and obtain
$\frac{1}{2}\lVert\rho^{\prime}_{n}-\psi_{n}\rVert_{1}<\frac{L\pi+2}{8L\pi}\epsilon\ .$ (113)
Therefore, the final bound Eq. (112) can be established. ∎
IV.2.2 Mixedness reduction subroutine
In this section, we discuss a subroutine that allows us to keep the mixedness parameter $x_{n}$ low in any given step during the QDP protocol, which is the critical assumption in Lemma 14 and Corollary 15. The challenge here is that the mixedness parameter $x_{n}^{\prime}$, obtained after an iteration of QDP Grover recursion, increases from $x_{n-1}<\frac{\epsilon}{L\pi}$, whilst for the next iteration, we again need a state with mixedness parameter $x_{n}<\frac{\epsilon}{L\pi}$. Hence, we cannot directly use the state $\rho_{n}^{\prime}=\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})$ as an input for the next iteration.
In order to reduce $x^{\prime}_{n}$ to $x_{n}<\frac{\epsilon}{L\pi}$, we utilize a probabilistic protocol introduced in Ref. [21], which we will refer to as interferential mixedness reduction (IMR). In Lemma 16, we present the IMR protocol along with the proof, using notation that explicitly incorporates the mixedness parameter. Afterwards, in Proposition 17, we analyze the success probability of the mixedness reduction subroutine, which arises from using multiple IMR rounds.
Lemma 16 (Interferential mixedness reduction (IMR) [21]).
Let $\rho=(1-x)\psi+x\frac{\mathmybb{1}_{\rm rel}}{2}$ be a state in $\mathcal{S}_{\rm rel}$ with mixedness parameter $x\in[0,1)$. Using a controlled swap operation on the two copies of the state, i.e. $\rho^{\otimes 2}$, a purer state $\rho^{\prime}=(1-x^{\prime})\psi+x^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}$ with mixedness parameter $x^{\prime}=\frac{2+x}{4-2x+x^{2}}x\leq x$ can be prepared with success probability $(1-\frac{x}{2}+\frac{x^{2}}{4})$. This procedure is oblivious, i.e. no knowledge of the state $\psi$ or the subspace $\mathcal{H}_{\rm rel}$ is required.
Proof.
For completeness we provide a self-contained proof for this two-qubit protocol with details on symmetric subspace projection using an ancilla qubit. The presented proof adapts to our notation the original derivation found in the paragraph around Eq. (14) of Ref. [21]. We remark that Ref. [21] also contains better, yet more complicated, protocols using multiple copies of $\rho^{\otimes M}$ simultaneously.
Following [21], any two copies of a rank-2 state $\rho\in\mathcal{S}_{\rm rel}$ can be written as
$\rho^{\otimes 2}=\frac{2x-x^{2}}{4}\Psi^{-}_{12}+\left(\frac{x^{2}}{4}\psi^{\perp}\otimes\psi^{\perp}+\frac{2x-x^{2}}{4}\Psi^{+}_{12}+\frac{(2-x)^{2}}{4}\psi\otimes\psi\right)\ ,$ (114)
where $\psi^{\perp}\in\mathcal{S}_{\rm rel}$ is a pure state orthonormal to $\psi$ and $\Psi^{\pm}_{12}=|{\Psi^{\pm}}\rangle\!\langle{\Psi^{\pm}}|_{12}$ are Bell states, such that $|{\Psi^{\pm}}\rangle_{12}=\frac{1}{\sqrt{2}}(|{\psi\psi^{\perp}}\rangle_{12}\pm|{\psi^{\perp}\psi}\rangle_{12})$. Subscripts $1$ and $2$ denote the space each copy of $\rho$ lives in. The main protocol is to apply the projector $\frac{\mathmybb{1}_{12}+S_{12}}{2}$, where $\mathmybb{1}_{12}$ is the identity on the total space of two copies and $S_{12}$ is a swap between the first and the second copies. Note that
$\frac{\mathmybb{1}_{12}+S_{12}}{2}=\psi^{\perp}\otimes\psi^{\perp}+\Psi^{+}_{12}+\psi\otimes\psi\ ,$ (115)
after which $\Psi^{-}_{12}$ component will vanish from the state $\rho^{\otimes 2}$. This operation can be implemented by a linear combination of unitaries utilizing an additional ancilla qubit $|{0}\rangle_{A}$ [33], since both $\mathmybb{1}_{12}$ and $S_{12}$ are unitaries. To a state $|{0}\rangle\!\langle{0}|_{A}\otimes\rho^{\otimes 2}$ appended with ancilla, apply a unitary
$=\left(|{0}\rangle\!\langle{0}|_{A}+|{1}\rangle\!\langle{1}|_{A}\right)\otimes\frac{\mathmybb{1}_{12}+S_{12}}{2}+\left(|{1}\rangle\!\langle{0}|_{A}+|{0}\rangle\!\langle{1}|_{A}\right)\otimes\frac{\mathmybb{1}_{12}-S_{12}}{2}\ ,$ (116)
$+\left[|{1}\rangle\!\langle{0}|_{A}\otimes\left(\frac{\mathmybb{1}_{12}-S_{12}}{2}\right)\rho^{\otimes 2}\left(\frac{\mathmybb{1}_{12}+S_{12}}{2}\right)+\mathrm{h.c.}\right]\ .$ (117)
$\begin{cases}\frac{x^{2}}{4}\psi^{\perp}\otimes\psi^{\perp}+\frac{2x-x^{2}}{4}\Psi^{+}_{12}+\frac{(2-x)^{2}}{4}\psi\otimes\psi\ ,\quad&\text{if$|{0}\rangle_{A}$ is measured,}\\ \frac{2x-x^{2}}{4}\Psi^{-}_{12}\ ,\quad&\text{if $|{1}\rangle_{A}$ is measured,}\end{cases}$ (118)
with probability given by their traces, $1-\frac{x}{2}+\frac{x^{2}}{4}$ and $\frac{2x-x^{2}}{4}$, respectively. For our purpose, the protocol is successful when $|{0}\rangle_{A}$ is measured, in which case the normalized density matrix
$\Omega_{12}=\frac{4}{4-2x+x^{2}}\left(\frac{x^{2}}{4}\psi^{\perp}\otimes\psi^{\perp}+\frac{2x-x^{2}}{4}\Psi^{+}_{12}+\frac{(2-x)^{2}}{4}\psi\otimes\psi\right)$ (119)
remains. The reduced state from $\Omega_{12}$ for both subsystems are the same and given as
$\rho^{\prime}\coloneq\operatorname{Tr}_{1}\left[\Omega_{12}\right]=\operatorname{Tr}_{2}\left[\Omega_{12}\right]=(1-x^{\prime})\psi+x^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2},$ (120)
with $x^{\prime}=\frac{2+x}{4-2x+x^{2}}x<x$. The correlation between two subsystems in $\Omega_{12}$ might play a destructive role in our algorithm. Hence, we discard one of the copies and take only a single reduced state $\rho^{\prime}$. ∎
When $x\ll 1$, the success probability is close to $1$ and $x^{\prime}=\frac{x}{2}+\mathcal{O}(x^{2})$, i.e. we efficiently reduce the mixedness by just consuming copies of the same state, without changing the associated pure state $\psi$. Another potential merit of running IMR protocol is that we can estimate $x$ and thus $x^{\prime}$ by the success probability, which can be used to improve the process in Lemma 15 by modifying angles $\{s_{l}\}$ in ^$\hat{\mathcal{E}}^{(\rho)}_{s_{l}}$ according to the estimated value of mixedness parameter. Nevertheless, for simplicity of analysis, we restrict ourselves to the case where $x$ is completely unknown apart from its upper bound.
Lemma 17 (QDP mixedness reduction subroutine).
Let $\rho=(1-x)\psi+x\frac{\mathmybb{1}_{\rm rel}}{2}$ be a state in $\mathcal{S}_{\rm rel}$ with mixedness parameter $x\in[0,\varepsilon]$ for some $0\leq\varepsilon<\frac{1}{2}$. Given any maximum tolerable probability of failure $q_{\mathrm{th}}\geq 0$, one can choose any parameter $\mathsf{g}\in\mathbb{R}$ and prepare $M$ copies of $\rho^{\prime}=(1-x^{\prime})\psi+x^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}$, such that:
  1. 1.
    $x^{\prime}\in[0,\frac{\varepsilon}{\mathsf{g}}]$,
  2. 2.
  3. 3.
    $c=1-\frac{\varepsilon}{2}-M^{-\frac{1}{2}}\sqrt{\log\left(\frac{R}{q_{\rm th}}\right)}\ ,$ (121)
  4. 4.
Note that $M$ must be sufficiently large to guarantee $c>0$.
Proof.
To prove this lemma, we analyze $R$ sequential rounds of the IMR protocol, with initial and final states $\rho$ and $\rho^{\prime}$. Let us denote the intermediate states generated via the IMR protocols as $\{\chi_{j}\}_{j=0}^{R}$, where $\chi_{0}=\rho$ and $\chi_{R}=\rho^{\prime}$; and let $\{y_{j}\}_{j=0}^{R}$ be the corresponding mixedness parameter for each $\chi_{j}$ with $y_{0}=x$ and $y_{R}=x^{\prime}$. Furthermore, let us denote $M_{j}$ to be the number of copies of an intermediate state $\chi_{j}$ generated; hence, we obtain $M_{R}=M$ copies of the desired state $\rho^{\prime}$ at the end, and aim to show that $M_{0}=M\times(\frac{2}{c})^{R}$ suffices. From Lemma 16, the mixedness parameter after $j$ rounds of the IMR protocol becomes
$y_{j}=\frac{2+y_{j-1}}{4-2y_{j-1}+y_{j-1}^{2}}y_{j-1}\leq\left(\frac{1}{2}+\dfrac{3y_{j-1}}{5}\right)y_{j-1}\ ,$ (122)
where the last inequality is obtained by comparing derivatives of the LHS and the RHS. The number of rounds $R$ is determined by the rate of mixedness reduction we desire. The final reduction $\frac{x^{\prime}}{x}=\frac{y_{R}}{y_{0}}\leq\mathsf{g}^{-1}$ is achievable if
$\frac{y_{R}}{y_{0}}=\prod_{j=1}^{R}\frac{y_{j}}{y_{j-1}}\leq\prod_{j=1}^{R}\left(\frac{1}{2}+\dfrac{3y_{j-1}}{5}\right)\leq\left(\frac{1}{2}+\frac{3\varepsilon}{5}\right)^{R}\leq\mathsf{g}^{-1}\ ,$ (123)
since $y_{j}\leq\varepsilon$ for all $j$. Therefore, $R=\mathcal{O}\left(\log(\mathsf{g})\right)$ is sufficient as claimed – note that $\log\left(\frac{1}{2}+\frac{3\varepsilon}{5}\right)$ to be $\mathcal{O}(1)$ since $\varepsilon<\frac{1}{2}$. The parameter $c$ in the statement of the lemma can be interpreted as the ‘survival rate’ after each round of the IMR protocol. More specifically, the $j$-th round of the protocol is successful if at least $\frac{c}{2}M_{j-1}$ output states $\chi_{j}$ is prepared from $M_{j-1}$ copies of $\chi_{j-1}$. If successful, we discard all the surplus output copies and set $M_{j}=\frac{c}{2}M_{j-1}$. If not, we declare that the whole subroutine has failed. Hence, if the entire subroutine succeeds, the final number of copies $M=M_{R}$ is related to the initial number of copies as $M_{0}=(2/c)^{R}M$. Finally, we estimate the success probability $q_{\rm succ}$ given $R,M,c,x$ and require it to be lower bounded with $(1-q_{\rm th})$. The failure probability of each round will be bounded using Hoeffding’s inequality. From Lemma 16, the success probability of preparing one copy of $\chi_{j}$ from a pair $\chi_{j-1}^{\otimes 2}$ is
$p(y_{j-1})=1-\frac{y_{j-1}}{2}+\frac{y_{j-1}^{2}}{4}\geq 1-\frac{y_{j-1}}{2}\ .$ (124)
At each round, $\frac{M_{j-1}}{2}$ independent trials of this IMR protocol are conducted. The probability of more than $c\frac{M_{j-1}}{2}$ attempts succeeds, i.e. the success probability of a round, is
$q_{r}^{(j)}=1-F\left(\left\lceil\frac{cM_{j-1}}{2}-1\right\rceil;\frac{M_{j-1}}{2},p(y_{j-1})\right)\geq 1-F\left(\frac{cM_{j-1}}{2};\frac{M_{j-1}}{2},1-\frac{y_{j-1}}{2}\right)\ ,$ (125)
$F(cn;n,p)=\sum_{i=0}^{\lfloor cn\rfloor}\binom{n}{i}p^{i}(1-p)^{n-i}\ .$ (126)
$q_{r}^{(j)}\geq 1-e^{-M_{j}\left(1-\frac{y_{j-1}}{2}-c\right)^{2}}\ .$ (127)
The success probability of an entire subroutine can then be bounded as
$q_{\rm succ}$ $=\prod_{j=1}^{R}q_{r}^{(j)}\geq\prod_{j=1}^{R}\left(1-e^{-M_{j}\left(1-\frac{y_{j-1}}{2}-c\right)^{2}}\right)>1-\sum_{j=1}^{R}e^{-M_{j}\left(1-\frac{y_{j-1}}{2}-c\right)^{2}}\ ,$ (128)
where the last inequality uses $\prod_{j}(1-q_{j})>1-\sum_{j}q_{j}$ that holds when $q_{j}\in(0,1)$ for all $j$. Moreover, by recalling $\varepsilon\geq y_{j-1}$ and $M\leq M_{j}$ for all $j$, the final bound
$q_{\rm succ}>1-Re^{-M\left(1-\frac{\varepsilon}{2}-c\right)^{2}}$ (129)
is obtained. The prescribed failure threshold $q_{\rm th}$ holds when $e^{-M\left(1-\frac{\varepsilon}{2}-c\right)^{2}}=\frac{q_{\rm th}}{R}$ , or equivalently, when $M$ is sufficently large and Eq. (121) is true. ∎
Note that the upper bound for $c$ in Eq. (121) increases monotonically as $M$ increases. Hence, the process is more efficient in earlier iterations of the Grover recursion, when $M$, the number of output states required, is larger. The worst bound is achieved at the last iteration. Ideally, we want $M=1$ at the last iteration, since the goal of the Grover search algorithm is to prepare a single copy of the state $\rho_{N}$ close to the target state. However, the threshold failure probability $q_{\rm th}$ cannot be arbitrarily small in such case, even if $c=0$. To guarantee an arbitrarily small failure probability, we prepare redundant copies of $\rho_{N}$ and denote the number of these copies as $M_{r}$. If we allow the redundancy $M_{r}>1$, it is also possible to fix $c$ to be a constant. From now on, let us fix
$c=1-\frac{\varepsilon}{2}-M_{r}^{-\frac{1}{2}}\sqrt{\log\left(\frac{R}{q_{\rm th}}\right)}=\frac{2}{3}\ ,$ (130)
which can be achieved by modulating $M_{r}$ since $1-\frac{\varepsilon}{2}>\frac{3}{4}$.
In addition, we find $R=\mathcal{O}(\log(\mathsf{g}))$ by setting $\mathsf{g}$. In Eq. (111), the mixedness of $\rho^{\prime}_{n}$ is bounded by $\frac{L\pi+2}{4L\pi}\epsilon$, while we want a state $\rho_{n}$ whose mixedness parameter is reduced to $\frac{\epsilon}{L\pi}$. Hence, the reduction by the factor $\mathsf{g}=\frac{L\pi+2}{4}$ is needed and therefore $R=\mathcal{O}(\log(L))$. Combining all this, Eq. (130) gives the redundancy
$M_{r}=\left(\frac{1}{3}-\frac{\varepsilon}{2}\right)^{-2}\log\left(\frac{\log(L)}{q_{\rm th}}\right)=\mathcal{O}\left(\log\left(\frac{1}{q_{\rm th}}\right)\right)\ ,$ (131)
assuming that the subroutine succeeds almost surely, i.e. $q_{\rm th}\lesssim\frac{1}{\log(L)}$.
We assumed that $\varepsilon<\frac{1}{2}$ in Lemma 17, but this is always true for our algorithm. $\varepsilon$ is given as the mixedness parameter of $\rho^{\prime}_{n}$ after a DME implementation Lemma 14, which is bounded by $\frac{L\pi+2}{4L\pi}\epsilon<0.41\epsilon<\frac{1}{2}$, since $\epsilon<1$.
@@ IV.3 QDP Grover search with linear circuit depth and exponential circuit width
In this subsection, we combine the results from Section IV.2 to prove a lemma describing a full QDP nested fixed-point Grover search algorithm.
@@
  • $\epsilon\in(0,1)$ be the threshold implementation error parameter,
  • $p_{\mathrm{th}}\lesssim\frac{N}{\log(L)}$ be the threshold failure probability of the algorithm.
@@
    $\delta_{n}=\operatorname{sech}\left((2L+1)\operatorname{arcsech}\left(\delta_{n-1}\right)\right)+\epsilon,$ (132)
@@
  • a number of initial state $\psi_{0}$ scaling as $L^{\mathcal{O}(N)}\epsilon^{-N}\log(p_{\rm th}^{-1})$.
@@
The algorithm consists of $N$ iterations of the recursion step, where each iteration involves two subroutines: a DME implementation in Lemma 14 followed by a mixedness reduction subroutine in Proposition 17.
@@
    $\rho_{n-1}=(1-x_{n-1})\psi_{n-1}+x_{n-1}\frac{\mathmybb{1}_{\rm rel}}{2}$ (133)
    is assumed to satisfy $x_{n-1}<\frac{\epsilon}{L\pi}$ and $\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\leq\delta_{n-1}$, while the output state
    $\rho_{n}^{\prime}=\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})=(1-x_{n}^{\prime})\psi_{n}+x_{n}^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}\ ,$ (134)
    satisfies $x^{\prime}_{n}<\frac{L\pi+2}{4L\pi}\epsilon$ from Eq. (111), and $\lVert\tau-\psi_{n}\rVert_{1}\leq\delta_{n}$ by Corollary 15. For each channel $\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}$ implementation, $\mathcal{O}(L^{2}\epsilon^{-1})$ DME queries are made and consequently $\mathcal{O}(L^{2}\epsilon^{-1})$ copies of $\rho_{n-1}$ are consumed. Plus, $L$ partial reflections $\hat{\mathbf{E}}^{(\tau)}_{s}$ around the target state are also applied. In sum, we need
    $\mathtt{I}_{n}=\mathcal{O}\left(\frac{L^{2}}{\epsilon}\right)\mathtt{O}_{n}$ (135)
@@
    At a mixedness reduction subroutine, described in Lemma 17, $\mathtt{O}_{n}$ copies of $\rho_{n}^{\prime}$, output from the DME implementation subroutine (Eq. (134)), are transformed into $\mathtt{I}_{n+1}$ copies of $\rho_{n}=(1-x_{n})\psi_{n}+x_{n}\frac{\mathmybb{1}_{\rm rel}}{2}$, such that $x_{n}<\frac{\epsilon}{L\pi}$, with a guarantee that the success probability $q_{\rm succ}$ is higher than $1-q_{\rm th}$ for some threshold value $q_{\rm th}$. The probability of $N$ such probabilistic subroutines to be all successful is
    $p_{\rm succ}=q_{\rm succ}^{N}\geq\left(1-q_{\rm th}\right)^{N}\geq 1-Nq_{\rm th}\ ,$ (136)
    where the second inequality follows from the Taylor expansion. Thus setting $q_{\rm th}=\frac{p_{\rm th}}{N}$ ensures that $p_{\rm succ}\geq 1-p_{\rm th}$ as required. Lemma 17 states $\mathtt{O}_{n}=\left(\frac{2}{c}\right)^{R}\mathtt{I}_{n+1}$ is needed, where $R$ is the number of IMR protocol rounds proportional to the $\log$ of rate of reduction $\frac{x_{n}^{\prime}}{x_{n}}$, while $c$ is the factor that decides how many transformed states survive after a round of the IMR protocol. As discussed in Eq. (130) and the following paragraph, we set $c=\frac{2}{3}$ and $R=\mathcal{O}(\log(L))$. Combining all,
    $\mathtt{O}_{n}=3^{\mathcal{O}\left(\log(L)\right)}\mathtt{I}_{n+1}=\mathrm{poly}(L)\mathtt{I}_{n+1}\ ,$ (137)
    and circuit depth $\mathcal{O}(R)=\mathcal{O}(\log(L))$ from $R$ repititions of interferential mixedness reduction protocols (Lemma 16) is added.
Finally, we account for the redundancy of the final number of copies $M_{r}$ in Eq. (131), introduced to guarantee high success probability. In other words, we set $I_{N}=M_{r}=\mathcal{O}(\log(Np_{\rm th}^{-1}))$. Then the initial number of copies needed
$I_{0}=L^{\mathcal{O}(N)}\epsilon^{-N}\log(p_{\rm th}^{-1})$ (138)
which is exponential in $N$. Furthermore, the circuit depth scales as
$\mathcal{O}\left(\log(L)+\frac{L^{2}}{\epsilon}\right)N=\mathcal{O}\left(\frac{NL^{2}}{\epsilon}\right)\ ,$ (139)

### v2

@@ IV.2.1 Distance bounds after a DME implementation
$\frac{1}{2}\lVert\rho^{\prime}_{n}-\tilde{\rho}_{n}\rVert_{1}\geq\left\lVert(x^{\prime}_{n}-x_{n-1})\psi_{n}-(x^{\prime}_{n}-x_{n-1})\frac{\mathmybb{1}_{\rm rel}}{2}\right\rVert_{1}=\lvert x^{\prime}_{n}-x_{n-1}\rvert.$ (S117)
@@
$\frac{1}{2}\lVert\rho^{\prime}_{n}-\tilde{\rho}_{n}\rVert_{1}\leq\frac{1}{2}\left\lVert\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\rho_{n-1})-\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\rho_{n-1})\right\rVert_{\rm 1}\leq\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}\right\rVert_{\rm Tr}<\frac{7L^{2}}{M}\ ,$ (S118)
from Eq. (S112). Therefore,
$0<x^{\prime}_{n}<x_{n-1}+\frac{7L^{2}}{M}<\frac{\epsilon}{4}\ ,$ (S119)
since we set $\frac{7L^{2}}{M}<\frac{L\pi-2}{4L\pi}\epsilon$ in Eq. (S116).
The mixedness factor $x^{\prime}_{n}$ can be reduced into $x_{n}\leq\frac{\epsilon}{2L\pi}$ using the IMR protocol Lemma 11, and we can continue the DME implementation without an accumulating error. Since we eventually change the factor $x^{\prime}_{n}$, it is more important to get a bound for $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}$ corresponding to the associated pure state, rather than $\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}$ as in Lemma 17.
@@
The resulting state $\rho^{\prime}_{n}=(1-2x_{n}^{\prime})\psi_{n}+x^{\prime}\mathmybb{1}_{\rm rel}$ of Lemma 17 has the associated pure state, such that
$\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}$ $\leq\delta_{n}\ ,$ (S120)
where $\delta_{n}$ is defined in Eq. (S99).
@@
The distance in Eq. (S120) follows the triangle inequality $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}\leq\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}+\frac{1}{2}\lVert\rho^{\prime}_{n}-\psi_{n}\rVert_{1}$. The first term in the RHS is already bounded by $\delta_{n}-\frac{\epsilon}{4}$ in Lemma 17. The second term $\frac{1}{2}\lVert\rho^{\prime}_{n}-\psi_{n}\rVert_{1}=x^{\prime}_{n}<\frac{\epsilon}{4}$ using Eq. (S119). ∎
@@ IV.3 QDP Grover search with linear circuit depth and exponential circuit width
In this subsection, we repeat the analysis in Theorem 12 for QDP nested fixed-point Grover search algorithm. This time, we can obtain scaling factors more precisely by combining the results from Section IV.2. As a result, we obtain the required number of initial state as a function of $L$, $\epsilon$, and $p_{\rm th}$ denoting the number of partial reflections of each recursion step, the threshold implementation error parameter, and the threshold failure probability.
@@
  • $\epsilon\in(0,\frac{2}{3})$ be the threshold implementation error parameter,
  • $p_{\mathrm{th}}\in(0,1)$ be the threshold failure probability of the algorithm.
@@
    $\delta_{n}=\operatorname{sech}\left((2L+1)\operatorname{arcsech}\left(\delta_{n-1}\right)\right)+\epsilon,$ (S121)
@@
  • a number of initial state $\psi_{0}$ scaling as $L^{(2+\Delta)N}\epsilon^{-N}\log(p_{\rm th}^{-1})$ with some positive constant $\Delta$.
@@
The algorithm consists of $N$ iterations of the recursion step, where each iteration involves two subroutines: a DME implementation in Lemma 17 followed by a mixedness reduction subroutine in Lemma 11. The proof is very similar to that of Theorem 12, except the parts where we explicitly consider the dependence on $L$.
@@
    $\rho_{n-1}=(1-2x_{n-1})\psi_{n-1}+x_{n-1}\mathmybb{1}_{\rm rel}$ (S122)
    is assumed to satisfy $x_{n-1}\leq\frac{\epsilon}{2L\pi}$ and $\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\leq\delta_{n-1}$, while the output state
    $\rho_{n}^{\prime}=\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\rho_{n-1})=(1-x_{n}^{\prime})\psi_{n}+x_{n}^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}\ ,$ (S123)
    satisfies $x^{\prime}_{n}<\frac{\epsilon}{4}$ from Eq. (S119), and $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}\leq\delta_{n}$ by Corollary 18. For each channel $\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}$ implementation, $\mathcal{O}(L^{2}\epsilon^{-1})$ DME queries are made and consequently $\mathcal{O}(L^{2}\epsilon^{-1})$ copies of $\rho_{n-1}$ are consumed. Plus, $L$ partial reflections $\hat{\mathbf{E}}^{[\tau]}_{s}$ around the target state are also applied. In sum, we need
    $\mathtt{I}_{n}=\mathcal{O}\left(\frac{L^{2}}{\epsilon}\right)\mathtt{O}_{n}$ (S124)
@@
    At a mixedness reduction subroutine, described in Lemma 11, $\mathtt{O}_{n}$ copies of $\rho_{n}^{\prime}$, output from the DME implementation subroutine (Eq. (S123)), are transformed into $\mathtt{I}_{n+1}$ copies of $\rho_{n}=(1-2x_{n})\psi_{n}+x_{n}\mathmybb{1}_{\rm rel}$, such that $x_{n}\leq\frac{\epsilon}{2L\pi}$, with a guarantee that the success probability $q_{\rm succ}$ is higher than $1-q_{\rm th}$ for some threshold value $q_{\rm th}$. Following Eqs. (S65) and (S66), we can set $q_{\rm th}=\frac{p_{\rm th}}{N}$ and $\mathtt{O}_{n}=(\frac{2}{c})^{R}\mathtt{I}_{n+1}$, where $R$ is the number of IMR protocol rounds proportional to the $\log$ of rate of reduction $\frac{x_{n}^{\prime}}{x_{n}}$, while $c$ is the factor that decides how many transformed states survive after a round of the IMR protocol. The rate of reduction $\frac{x_{n}^{\prime}}{x_{n}}=\frac{L\pi}{2}$, giving $R=\mathcal{O}(\log(\mathsf{g}))=\mathcal{O}(\log(L))$, while we can set $c=\frac{1}{2}$ from Eq. (S67). Combining all,
    $\mathtt{O}_{n}=4^{\mathcal{O}\left(\log(L)\right)}\mathtt{I}_{n+1}=\mathrm{poly}(L)\mathtt{I}_{n+1}\ ,$ (S125)
    and circuit depth $\mathcal{O}(R)=\mathcal{O}(\log(L))$ from $R$ repetitions of interferential mixedness reduction protocols (Lemma 10) is added.
Finally, we account for the redundancy of the final number of copies $I_{N+1}=\mathcal{O}(\log(Np_{\rm th}^{-1}))$ in Eq. (S67), introduced to guarantee high success probability and fix the survival rate $c$. Then the initial number of copies needed
$I_{1}=L^{(2+\Delta)N}\epsilon^{-N}\log(p_{\rm th}^{-1}),$ (S126)
where $\Delta$ is the additional exponent coming from the mixedness reduction rounds. The number of initial copies is exponential in $N$ and grows faster than $L^{2N}$ as a function of $L$. Furthermore, the circuit depth scales as
$\mathcal{O}\left(\log(L)+\frac{L^{2}}{\epsilon}\right)N=\mathcal{O}\left(\frac{NL^{2}}{\epsilon}\right)\ ,$ (S127)